OpenStudy (anonymous):

6t+5 3t+7 ___________ = ________ 4t-3 2t-7

OpenStudy (anonymous):

\[\frac{ A }{B } = \frac{ C }{ D } \rightarrow \frac{ AD }{ BC } = 1\]

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):